# Aligning Machine and Human Visual Representations across Abstraction Levels

### v2

@@ Abstract
Deep neural networks have achieved success across a wide range of applications, including as models of human behavior in vision tasks. However, neural network training and human learning differ in fundamental ways, and neural networks often fail to generalize as robustly as humans do, raising questions regarding the similarity of their underlying representations. What is missing for modern learning systems to exhibit more human-like behavior? We highlight a key misalignment between vision models and humans: whereas human conceptual knowledge is hierarchically organized from fine- to coarse-scale distinctions, model representations do not accurately capture all these levels of abstraction. To address this misalignment, we first train a teacher model to imitate human judgments, then transfer human-like structure from its representations into pretrained state-of-the-art vision foundation models. These human-aligned models more accurately approximate human behavior and uncertainty across a wide range of similarity tasks, including a new dataset of human judgments spanning multiple levels of semantic abstractions. They also perform better on a diverse set of machine learning tasks, increasing generalization and out-of-distribution robustness. Thus, infusing neural networks with additional human knowledge yields a best-of-both-worlds representation that is both more consistent with human cognition and more practically useful, thus paving the way toward more robust, interpretable, and human-like artificial intelligence systems.
@@ 1 Introduction
Figure 1: A: An example of the triplet odd-one-out task where a human and a neural network model choose the same (top) and a different (bottom) odd-one-out image respectively. B: The different parts of the AligNet framework depicted from end to end. First, we develop a teacher model of human judgments using the THINGS dataset (first panel). Second, we apply this model to ImageNet and cluster its latent representations into semantically meaningful categories (second panel). This allows us to generate arbitrarily many similarity judgments. Third, the obtained human-like similarity structure information is distilled into a student vision foundation model using a novel loss function (third panel). C: Representative human alignment (top panel) and machine learning downstream (bottom panel) results show significant performance improvements of the aligned over the non-aligned version of the ViT-B classifier. D: 2D latent space projection for visualizing the change in the representations after alignment. While the representations of a standard ViT-B classifier model are unstructured and categories overlap, after alignment the representations are grouped into meaningful categories.
@@
While deep learning has recently driven rapid progress in areas of artificial intelligence such as natural language processing [LeCun et al., 2015; Brown et al., 2020] and computer vision [Radford et al., 2021; Dosovitskiy et al., 2020; Zhai et al., 2023; Kirillov et al., 2023], even the best of these systems often fail in ways that humans would not [Lapuschkin et al., 2016; Geirhos et al., 2018; Lapuschkin et al., 2019; Geirhos et al., 2020; Hermann et al., 2019; Thrush et al., 2022; Bowers et al., 2022; Muttenthaler et al., 2023b; Hermann et al., 2023]. These failures have led to renewals [Lake et al., 2017; Bowers et al., 2022] of older arguments that neural networks lack essential ingredients of human intelligence [Fodor and Pylyshyn, 1988; Marcus, 1998; McClelland et al., 2010; Holyoak and Hummel, 2014; Tenenbaum et al., 2011]. How can we build systems that have more human-like behavior?
@@
Human perception is robust to changes in the environment and generalizes across different visual settings [Lake et al., 2015, 2017; Roads and Mozer, 2017; Geirhos et al., 2018; Lake and Baroni, 2023]. However, model performance declines—often drastically—if the data distribution between training and test sets is shifted [e.g., Sugiyama et al., 2007; Sugiyama and Kawanabe, 2012; Geirhos et al., 2020; Hendrycks et al., 2021a]. This lack of robustness in the representations of vision models poses a challenge for many downstream applications that require the ability to generalize [e.g., Lapuschkin et al., 2019; Pooch et al., 2019; Geirhos et al., 2020]. In addition, humans tend to be well-calibrated—for example, when they are asked to judge visual similarity [Roads and Mozer, 2017]—that is, humans’ (un)certainty tends to correlate with their (in)accuracy. AI systems, on the other hand, are often overconfident and show high certainty [Guo et al., 2017] even when their predictions are incorrect [Minderer et al., 2021]. Thus, many differences remain to be reconciled before we can ultimately achieve human-like artificial intelligence.
@@
One key step to assessing—and addressing—this misalignment is to be able to systematically and reliably compare human and machine judgments, e.g., of similarity or categorical membership. Such comparisons are daunting, however, because of the substantial difference in data availability between the two sources. Human behavioral data is scarce because collecting representative samples is challenging and expensive [Cichy et al., 2019; King et al., 2019; Hebart et al., 2020, 2023]. The cost of getting (human) similarity judgments scales with the square of the number of images. For this reason, the largest human similarity-judgment datasets to date are several orders of magnitudes smaller than the datasets used to pretrain deep learning models. To build more accurate models of human perception, we propose a novel framework that enables the simulation of arbitrarily large (albeit imperfect) human-like similarity judgment datasets for alignment via a teacher model to render decisions of large neural network models more human-like.
@@
We begin by using an affine transformation to align neural network model representations with human semantic judgments in triplet odd-one-out tasks [Muttenthaler et al., 2023a]. In doing so, we regularize the alignment to preserve the local representation structure of the models, and additionally leverage uncertainty measures over human responses to improve model calibration. This enables us to define a transformation toward more human-like representations of a state-of-the-art image/text vision foundation model (VFM). For the remainder of this paper, we use this aligned version of the VFM as a surrogate model for generating human similarity judgments. This approach is loosely analogous to the use of surrogate models for pseudo-labelling in semi-supervised classification learning [e.g. Lee, 2013]. With our surrogate model, we hypothesize that we can create a larger training dataset with multiple levels of conceptual granularity that will serve to provide cross-level constraints on human alignment.
@@
To do so, we group the representations of the surrogate model into meaningful superordinate categories which we use to sample semantically meaningful triplets. Subsequently, we obtain odd-one-out responses from our surrogate model for all the triplets that we generate. This results in a large dataset of human-like triplet odd-one-out responses. We call this dataset AligNet. To transfer this carefully curated human-like similarity structure into a pretrained neural network foundation model, we introduce a new Kullback-Leibler divergence based objective function that facilitates the learning process. The different steps of the AligNet framework are visualized in Fig. 1B. While we use the ImageNet dataset [Russakovsky et al., 2014] because it is publicly available and widely used in the cognitive science and computer vision communities, our framework allows for different versions of AligNet to be created based on other image datasets, irrespective of whether the datasets are labelled.
@@
In addition, AligNet yields a suite of open-source vision models that encode hierarchical human knowledge through soft-alignment [^1] Models will be open-sourced upon acceptance of the manuscript for publication.. Its general significance stems from enabling the sciences, medicine, and industries to use more human-like and robust vision models for any downstream application. The use of the ImageNet data to create AligNet gives a publicly available image dataset that can be employed by any researcher or practitioner without restriction.
@@ 2.1 Machine learning for alignment
To build models with more human-like behavior it is necessary to inject additional supervision about human behavior into the representations of foundation models. One way to achieve this is via an objective function that explicitly forces a model’s representation to inherit a human concept space [Muttenthaler et al., 2023a, b; Fu et al., 2023]. Previous work has shown that the objective function in addition to the (pre-)training data is the most central variable in the pursuit of human-like vision models [Muttenthaler et al., 2023a; Conwell et al., 2023].
@@ 2.2 Toward human-like systems via a surrogate approach
Triplet odd-one-out task. Models fine-tuned on AligNet achieved odd-one-out accuracies of $60.45\%$ (ViT-B) -- $62.54\%$ (SigLIP ViT-B) of which $62.54\%$ is the best alignment score of a neural network model on the THINGS dataset to date [c.f. Muttenthaler et al., 2023a, b]. In contrast, the base model representations achieved odd-one-out accuracies of $34.87\%$ (ViT-B) -- $51.21\%$ (DINOv2 ViT-B). The odd-one-out accuracy of ViT-B increased substantially from $34.87\%$ to $60.45\%$. This is a relative increase in performance of $73.35\%$. All performance increases are statistically significant at $\alpha=0.05$.
@@ 2.2.2 Alignment at multiple levels of abstraction
Local fine-grained. The performance of most base models (except for DINOv1 ViT-B) did not strongly correspond to human responses for fine-grained semantics either and, thus, model performances were far from the human noise ceiling of $65.92\%$. Prior to fine-tuning, models achieved poor alignment scores of $46.04\%$ (ViT-B) -- $57.72\%$ (DINOv2 ViT-B), except for DINOv1 ViT-B which performed ($62.92\%$) significantly better than any other model in that setting (see Tab. 3 in the Supplementary Material). AligNet fine-tuning improved this mismatch to some degree but not as substantially as it did for the coarse-grained abstraction level. AligNet models achieved odd-one-out accuracies of $58.93$ (ViT-S) -- $62.92\%$ (DINOv1 ViT-B). This is equal to a relative increase in performance of $7.84\%$ (DINOv2 ViT-B) - $46.03\%$ (ViT-L) (see Fig. 2D, middle column).
@@ 2.3 Aligned models reflect the conceptual hierarchy
How do the model representations change after soft-alignment? In Fig. 3 we show that while the model representations are dissimilar before alignment (especially those trained on classification tasks), they become much more similar to each other after soft-alignment. (See Sec. 4.4 for details.)
@@
As an illustrative example, in a standard ViT-B classifier the representations of lizards are initially closer to those of some plants and fruits due their similarity in texture, color, or background conditions; after alignment, they are naturally more similar to the representations of other animals and more distant from those of other, unrelated superordinate categories (see Figure 1D). This organization will yield better generalization when lizards are anomalous among a category of fruits, or when a lizard image is used as one of the examples depicting a more abstract category like animals.
@@ 2.4 Alignment improves generalization and out-of-distribution robustness
One-shot classification. One-shot classification tasks are among the hardest scenarios for testing generalization: the objective is to classify images given only a single labelled example per class. In Fig. 4 we show one-shot performance of all models before and after AligNet finetuning on a ten image-classification datasets from a wide variety of domains, such as fine-grained birds classification [Birds; Welinder et al., 2010], land use classification from satellite imagery [UC Merced; Yang and Newsam, 2010], and colorectal cancer histology [Colon; Kather et al., 2016]. While for eight combinations of model and dataset, we find a small decrease in performance, the majority of cases (32) show an improvement, sometimes by a substantial margin (e.g. accuray of DINOv2 on Pets shows a 2.7 fold increase). Overall our alignment framework increased the generalization performance on these task significantly (p < 0.05). These results corroborate the hypothesis that a good coarse-level (re-)structuring of the representation space supports generalization in general and especially for tasks where little labelled data is available.
@@
Distribution shift. A long-standing problem of applying machine learning algorithms in practice is distribution shift and domain adaptation [cf. Sugiyama et al., 2007; Sugiyama and Kawanabe, 2012; Farahani et al., 2021]: as machine learning systems are deployed, the data they were originally trained on often differs in subtle ways from the data they are being applied to. Models tend to fail in unexpected ways when the data they should predict has a different distribution than the training data.
To measure if the global category structure induced by fine-tuning on AligNet helps to address such distributional shifts, we use the BREEDS benchmarks [Santurkar et al., 2020]. These were specifically designed explore generalization under input distributions shifts, by constructing datasets where the samples from training and test are sampled from different sub-populations. Fig. 4B shows the result of training a linear readout layer on the four BREEDS training sets, and evaluating them on the corresponding test-sets. AligNet fine-tuning consistently improves performance across all benchmarks and models types. The image classifier (ViT-B) benefits the most with a more than twofold increase for the “entity13” benchmark.
Model robustness. A further benefit of alignment is that it has a positive influence on the robustness of a model. Hendrycks et al. [2021b] collected a challenging dataset of natural images that are adversarial, in the sense that neural networks tend to consistently misclassify them, while they are obvious for humans. Again, we find that AligNet fine-tuning improves the accuracy across all VFMs that we investigate (cf. Fig. 4C); specifically, we observe improvements of up to $9.5$ percentage points, which corresponds to a $1.6$ fold improvement for the CapPa model.
Taken together, all of these results corroborate that aligning human and neural network model representations helps to generalize better, transfer to new tasks and data, and enhances model robustness. In other words, alignment is helpful for practically improving deep learning.
@@ 3 Discussion
The differences between natural intelligence and the capabilities of neural networks are the subject of long-standing debates [Fodor and Pylyshyn, 1988; Lake et al., 2017]. Despite the dramatic progress in AI of the past decade, these discussions persist, because even the most successful deep learning systems seem to fail in non-human-like ways [Lake et al., 2015; Lapuschkin et al., 2019; Geirhos et al., 2020; Thrush et al., 2022; Muttenthaler et al., 2023a; Lake and Baroni, 2023; Puatruaucean et al., 2023].
In this work, we have highlighted—and addressed—a key deficiency in a broad class of vision foundation models: their representations do not adequately represent the multi-level conceptual structure of semantic knowledge that is natural to humans (see Sec. 2.2). We demonstrate this deficiency by collecting a new dataset of stratified human similarity judgments across multiple levels of abstraction—which we call Levels (see Sec. 2.2.2). To address this deficiency, we established a methodological framework for aligning deep learning models with human similarity judgments to build systems that exhibit more human-like behavior (see Sec. 2.1). By bootstrapping from a relatively small quantity of human data, we develop a surrogate neural network model of human object similarity judgments. Analogous to prior work on semi-supervised classification [e.g. Lee, 2013; Grandvalet and Bengio, 2004], we use this surrogate model to label a larger dataset for which we do not have human judgments. Specifically, we propose a cluster-based triplet-sampling technique for ImageNet images, which we then label with human-like odd-one-out responses from our surrogate model to create a large new synthetic dataset called AligNet.
@@
We remark that AI systems have been successfully adopted in many areas across the sciences and industries. However, these deployments lead to many practical [Paleyes et al., 2023] and conceptual [Amodei et al., 2016] concerns such as trustworthiness, biases, safety issues, and miscalibration. It is therefore increasingly important to identify the reasons why these systems occasionally fail (e.g., global misalignment, non-robustness, spurious correlations, or shortcut learning) and how to alleviate these failures. Our work advances the understanding of the deficiencies of a major class of current systems, and simultaneously shows a viable path for ameliorating these deficiencies by alignment with human judgments.
In order to most effectively build on this work, there are a number of limitations that could be addressed in future efforts. First, the neural network models we used neither account for context in similarity judgments nor higher-order relations. These factors could possibly be accounted for with more complex models, such as contemporary vision-language models [e.g. Team et al., 2023; OpenAI et al., 2023], with which performance and alignment across different contexts could be readily explored. Additionally, in this work we generally treat the human representation as if it were unique—stressing the intrinsic uncertainty across individuals. However, human representations may vary systematically across individuals, cultures, and so on. Exploring this dimension of variation in human judgments would be an important topic for future investigation. Finally, human judgment is full of flaws, intrinsic contradictions and discrepancies. Given these issues, perfect alignment to human performance may not always be desirable for a technical system. Thus, future work could explore how to best learn from human knowledge without adopting human imperfections.

### v3

@@ Abstract
Deep neural networks have achieved success across a wide range of applications, including as models of human behavior in vision tasks. However, neural network training and human learning differ in fundamental ways, and neural networks often fail to generalize as robustly as humans do, raising questions regarding the similarity of their underlying representations. What is missing for modern learning systems to exhibit more human-like behavior? We highlight a key misalignment between vision models and humans: whereas human conceptual knowledge is hierarchically organized from fine- to coarse-scale distinctions, model representations do not accurately capture all these levels of abstraction. To address this misalignment, we first train a teacher model to imitate human judgments, then transfer human-like structure from its representations to refine the representations of pretrained state-of-the-art vision foundation models via finetuning. These human-aligned models more accurately approximate human behavior and uncertainty across a wide range of similarity tasks, including a new dataset of human judgments spanning multiple levels of semantic abstractions. They also perform better on a diverse set of machine learning tasks, increasing generalization and out-of-distribution robustness. Thus, infusing neural networks with additional human knowledge yields a best-of-both-worlds representation that is both more consistent with human cognition and more practically useful, thus paving the way toward more robust, interpretable, and human-like artificial intelligence systems.
@@ 1 Introduction
Figure 1: a: An example of the triplet odd-one-out task where a human and a neural network model choose the same (top) and a different (bottom) odd-one-out image respectively. b: The different parts of the AligNet framework depicted from end to end. First, we develop a teacher model of human judgments using the THINGS dataset (first panel). Second, we apply this model to ImageNet and cluster its latent representations into semantically meaningful categories (second panel). This allows us to generate arbitrarily many similarity judgments. Third, the obtained human-like similarity structure information is distilled into a student vision foundation model using a novel loss function (third panel). c: Representative human alignment (top panel) and machine learning downstream (bottom panel) results show significant performance improvements of the aligned over the non-aligned version of the ViT-B classifier (up to 123% of relative improvement for ML downstream transfer). d: 2D latent space projection for visualizing the change in the representations after alignment. While the representations of a standard ViT-B classifier model are unstructured and categories overlap, after alignment the representations are grouped into meaningful categories.
@@
While deep learning has recently driven rapid progress in areas of artificial intelligence such as natural language processing [LeCun et al., 2015; Brown et al., 2020] and computer vision [Radford et al., 2021; Dosovitskiy et al., 2021; Zhai et al., 2023; Kirillov et al., 2023], even the best of these systems often fail in ways that humans would not [Lapuschkin et al., 2016; Geirhos et al., 2018; Lapuschkin et al., 2019; Geirhos et al., 2020; Hermann et al., 2020; Thrush et al., 2022; Bowers et al., 2022; Muttenthaler et al., 2023b; Hermann et al., 2023]. These failures have led to renewals [Lake et al., 2017; Bowers et al., 2022] of older arguments that neural networks lack essential ingredients of human intelligence [Fodor and Pylyshyn, 1988; Marcus, 1998; McClelland et al., 2010; Holyoak and Hummel, 2014; Tenenbaum et al., 2011]. How can we build systems that have more human-like behavior?
@@
Human perception is robust to changes in the environment and generalizes across different visual settings [Lake et al., 2015, 2017; Roads and Mozer, 2017; Geirhos et al., 2018; Lake and Baroni, 2023]. However, model performance declines—often drastically—if the data distribution between training and test sets is shifted [e.g., Sugiyama et al., 2007; Sugiyama and Kawanabe, 2012; Geirhos et al., 2020; Hendrycks et al., 2021a]. This lack of robustness in the representations of vision models poses a challenge for many downstream applications that require the ability to generalize [e.g., Lapuschkin et al., 2019; Pooch et al., 2020; Geirhos et al., 2020]. In addition, humans tend to be well-calibrated—for example, when they are asked to judge visual similarity [Roads and Mozer, 2017]—that is, humans’ (un)certainty tends to correlate with their (in)accuracy. AI systems, on the other hand, are often overconfident and show high certainty [Guo et al., 2017] even when their predictions are incorrect [Minderer et al., 2021]. Thus, many differences remain to be reconciled before we can ultimately achieve human-like artificial intelligence.
@@
One key step to assessing—and addressing—this misalignment is to be able to systematically and reliably compare human and machine judgments, e.g., of similarity or categorical membership. Such comparisons are daunting, however, because of the substantial difference in data availability between the two sources. Human behavioral data is scarce because collecting representative samples is challenging and expensive [Cichy et al., 2019; King et al., 2019; Hebart et al., 2020, 2023]. The cost of getting (human) similarity judgments scales with the square of the number of images. For this reason, the largest human similarity-judgment datasets to date are several orders of magnitudes smaller than the datasets used to pretrain deep learning models. To build more accurate models of human perception, we propose a novel framework that enables the simulation of arbitrarily large (albeit imperfect) human-like similarity judgment datasets for alignment via a teacher model. We use the created datasets to fine-tune large pretrained neural network models to yield more human-like representations and behavior. This approach allows substantially improving human-model alignment of a broad range of vision foundation models, without repeating the entire pretraining process.
@@
We begin by using an affine transformation to align neural network model representations with human semantic judgments in triplet odd-one-out tasks [Muttenthaler et al., 2023a]. In doing so, we regularize the alignment to preserve the local representation structure of the models, and additionally leverage uncertainty measures over human responses to improve model calibration. This enables us to define a transformation toward more human-like representations of a state-of-the-art image/text vision foundation model (VFM). For the remainder of this paper, we use this aligned version of the VFM as a surrogate model for generating human similarity judgments. This approach is loosely analogous to the use of surrogate models for pseudo-labelling in semi-supervised classification learning [e.g. Lee et al., 2013]. With our surrogate model, we hypothesize that we can create a larger training dataset with multiple levels of conceptual granularity that will serve to provide cross-level constraints on human alignment.
@@
To do so, we group the representations of the surrogate model into meaningful superordinate categories which we use to sample semantically meaningful triplets. Subsequently, we obtain odd-one-out responses from our surrogate model for all the triplets that we generate. This results in a large dataset of human-like triplet odd-one-out responses. We call this dataset AligNet. To transfer this carefully curated human-like similarity structure into a pretrained neural network foundation model, we introduce a new Kullback-Leibler divergence based objective function that facilitates the learning process. The different steps of the AligNet framework are visualized in Fig. 1c. While we use the ImageNet dataset [Russakovsky et al., 2015] because it is publicly available and widely used in the cognitive science and computer vision communities, our framework allows for different versions of AligNet to be created based on other image datasets, irrespective of whether the datasets are labelled.
@@
In addition, AligNet yields a suite of open-source adaptations of vision foundation models that encode hierarchical human knowledge through soft-alignment [^1] Models will be open-sourced upon acceptance of the manuscript for publication.. Its general significance stems from enabling the sciences, medicine, and industries to use more human-like and robust vision models for any downstream application. The use of the ImageNet data to create AligNet gives a publicly available image dataset that can be employed by any researcher or practitioner without restriction.
@@ 2.1 Machine learning for alignment
To build models with more human-like behavior it is necessary to inject additional supervision about human behavior into the representations of foundation models. One way to achieve this is via an objective function that explicitly forces a model’s representation to inherit a human concept space [Muttenthaler et al., 2023a, b; Fu et al., 2023]. Previous work has shown that the objective function in addition to the (pre-)training data is the most central variable in the pursuit of human-like vision models [Muttenthaler et al., 2023a; Conwell et al., 2022].
@@ 2.2 Toward human-like systems via a surrogate approach
Triplet odd-one-out task. Models fine-tuned on AligNet achieved odd-one-out accuracies of $60.45\%$ (ViT-B) -- $62.54\%$ (SigLIP ViT-B) of which $62.54\%$ is the best alignment score of a neural network model on the THINGS dataset to date [cf. Muttenthaler et al., 2023a, b]. In contrast, the base model representations achieved odd-one-out accuracies of $34.87\%$ (ViT-B) -- $51.21\%$ (DINOv2 ViT-B). The odd-one-out accuracy of ViT-B increased substantially from $34.87\%$ to $60.45\%$. This is a relative increase in performance of $73.35\%$. All performance increases are statistically significant at $\alpha=0.05$.
@@ 2.2.2 Alignment at multiple levels of abstraction
Local fine-grained. The performance of most base models (except for DINOv1 ViT-B) did not strongly correspond to human responses for fine-grained semantics either and, thus, model performances were far from the human noise ceiling of $65.92\%$. Prior to fine-tuning, models achieved poor alignment scores of $46.04\%$ (ViT-B) -- $57.72\%$ (DINOv2 ViT-B), except for DINOv1 ViT-B which performed ($62.92\%$) significantly better than any other model in that setting (see Tab. 3 in the Supplementary Material). AligNet fine-tuning improved this mismatch to some degree but not as substantially as it did for the coarse-grained abstraction level. AligNet models achieved odd-one-out accuracies of $58.93$ (ViT-S) -- $62.92\%$ (DINOv1 ViT-B). This is equal to a relative increase in performance of $7.84\%$ (DINOv2 ViT-B) - $46.03\%$ (ViT-L) (see Fig. 2d, center column).
@@ 2.3 Aligned models reflect the conceptual hierarchy
How do the model representations change after soft-alignment? In Fig. 3 we show that while the model representations are dissimilar before alignment (especially those trained on classification tasks), they become much more similar to each other after soft-alignment (see Sec. 4.4 for details).
@@
As an illustrative example, in a standard ViT-B classifier the representations of lizards are initially closer to those of some plants and fruits due their similarity in texture, color, or background conditions; after alignment, they are naturally more similar to the representations of other animals and more distant from those of other, unrelated superordinate categories (see Fig. 1d). This organization will yield better generalization when lizards are anomalous among a category of fruits, or when a lizard image is used as one of the examples depicting a more abstract category like animals.
@@ 2.4 Alignment improves generalization and out-of-distribution robustness
One-shot classification. One-shot classification tasks are among the hardest scenarios for testing generalization: the objective is to classify images given only a single labelled example per class. In Fig. 4 we show one-shot performance of all models before and after soft-alignment on ten image-classification datasets from a wide range of domains, such as fine-grained bird types classification [Birds; Wah et al., 2011], land use classification from satellite imagery [UC Merced; Yang and Newsam, 2010], and colorectal cancer histology [Colon; Kather et al., 2016]. While for eight combinations of models and datasets, we find a small decrease in performance, the majority of cases (32) show an improvement, sometimes by a substantial margin (e.g., the accuracy of DINOv2 on the Pets dataset shows a 2.7 fold increase). Overall our alignment framework increased the generalization performance on these task significantly (p < 0.05). These results corroborate the hypothesis that a good coarse-level (re-)structuring of the representation space supports generalization in general and especially for tasks where little labelled data is available. In Appx. B.2.1 we show additional few-shot results for more tasks and models, and demonstrate that our method can be used in conjunction with other approaches that were specifically designed for improving the performance of pretrained foundation models in few-shot learning tasks. Often a caveat with those methods is that their application is restricted to image/text models because they require text representations [e.g., Zhang et al., 2022]. In contrast, soft-alignment can be applied to a model irrespective of how it was pretained.
@@
Distribution shift. A long-standing problem of applying machine learning algorithms in practice is distribution shift and domain adaptation [cf., Sugiyama et al., 2007; Sugiyama and Kawanabe, 2012; Farahani et al., 2021]: as machine learning systems are deployed, the data they were originally trained on often differs in subtle ways from the data they are being applied to. Models tend to fail in unexpected ways when the distribution of the data they should predict differs from that of the training data.
To measure whether the global category structure induced by fine-tuning on AligNet helps address these distributional shifts, we use the BREEDS benchmarks [Santurkar et al., 2020]. These were specifically designed explore generalization under input distribution shifts, by constructing datasets where the samples from training and test are sampled from different subpopulations. Fig. 4b shows the result of training a linear readout layer on the four BREEDS training sets, and evaluating them on the corresponding test sets. AligNet fine-tuning consistently improves performance across all benchmarks and models types. The image classifier (ViT-B) benefits the most with a more than twofold increase for the “entity13” benchmark.
Model robustness.
A further benefit of alignment is that it has a positive influence on the robustness of a model. Hendrycks et al. [2021b] collected a challenging dataset of natural images that are adversarial, in the sense that neural networks tend to consistently misclassify them, while they are obvious for humans. Again, we find that AligNet fine-tuning improves the accuracy across all VFMs that we investigate (cf. Fig. 4c); specifically, we observe improvements of up to $9.5$ percentage points, which corresponds to a $1.6$ fold improvement for the CapPa model.
Together, all of these results corroborate that aligning human and neural network model representations helps to generalize better, transfer to new tasks and data, and enhances model robustness. In other words, alignment is helpful to practically improve deep learning.
@@ 3 Discussion
The differences between natural intelligence and the capabilities of neural networks are the subject of long-standing debates [Fodor and Pylyshyn, 1988; Lake et al., 2017]. Despite the dramatic progress in AI of the past decade, these discussions persist, because even the most successful deep learning systems seem to fail in non-human-like ways [Lake et al., 2015; Lapuschkin et al., 2019; Geirhos et al., 2020; Thrush et al., 2022; Muttenthaler et al., 2023a; Lake and Baroni, 2023; Patraucean et al., 2023].
In this work, we have highlighted—and addressed—a key deficiency in a broad class of vision foundation models: their representations do not adequately represent the multi-level conceptual structure of semantic knowledge that is natural to humans (see Sec. 2.2). We demonstrate this deficiency by collecting a new dataset of stratified human similarity judgments across multiple levels of abstraction—which we call Levels (see Sec. 2.2.2). To address this deficiency, we established a methodological framework for aligning deep learning models with human similarity judgments to build systems that exhibit more human-like behavior (see Sec. 2.1). By bootstrapping from a relatively small quantity of human data, we develop a surrogate neural network model of human object similarity judgments. Analogous to prior work on semi-supervised classification [e.g. Lee et al., 2013; Grandvalet and Bengio, 2004], we use this surrogate model to label a larger dataset for which we do not have human judgments. Specifically, we propose a cluster-based triplet-sampling technique for ImageNet images, which we then label with human-like odd-one-out responses from our surrogate model to create a large new synthetic dataset called AligNet.
@@
We remark that AI systems have been successfully adopted in many areas across the sciences and industries. However, these deployments lead to many practical [Paleyes et al., 2022] and conceptual [Amodei et al., 2016] concerns such as trustworthiness, biases, safety issues, and miscalibration. It is therefore increasingly important to identify the reasons why these systems occasionally fail (e.g., global misalignment, non-robustness, spurious correlations, or shortcut learning) and how to alleviate these failures. Our work advances the understanding of the deficiencies of a major class of current systems, and simultaneously shows a viable path for ameliorating these deficiencies by alignment with human judgments.
In order to most effectively build on this work, there are a number of limitations that could be addressed in future efforts. First, the neural network models we used neither account for context in similarity judgments nor higher-order relations. These factors could possibly be accounted for with more complex models, such as contemporary vision-language models [e.g. Team et al., 2023; OpenAI et al., 2023], with which performance and alignment across different contexts could be readily explored. In this work we generally treat the human representation as if it were common among all humans, modulo some intrinsic uncertainty across individuals. However, human representations may vary systematically across individuals, cultures, and so on. Exploring this dimension of variation in human judgments would be an important topic for future investigation. Finally, human judgment is full of flaws, intrinsic contradictions and discrepancies. Given these issues, perfect alignment to human performance may not always be desirable for a technical system. Thus, future work could explore how to best learn from human knowledge without adopting human imperfections.